27. Find det(A – 21) if A is the matrix given in Problem 4.-1 4. 2 0 - 11 3 2 -3 1
Answer #1

2 0 A= 1 3 2 -3 de- 032 o 2 0 D odo Ads = 2-1 0 --1 1 می) 2 - 3rd au (Adf) det 2-1 - 1 0 Id 2 1 3 -3rd :)] = 2nd M -) (2-0) [